# Multi-User Chat Assistant (MUCA): a Framework Using LLMs to Facilitate Group Conversations

### 10 Jan 2024 v1

@@ A.2.4 User Simulation
Figure 17 shows full prompting of input prompt templates where purple texts are represented as placeholders and will be replaced when generating the input prompt for inference. All prompting use the CoT style, for example, a virtual user ID is first determined for the next turn and then its speaking role is decided based on the user ID. Figure 18 shows two examples of MUS for different topics. Examples demonstrate that MUS is able to mimic different user’s behavior based on their speaking roles and utterance traits.
A.2.5 System Design
@@
The user interface (UI), designed with JavaScript, HTML, and CSS, is a static single-page web application which is responsible for managing user login and facilitating communication with the backend server. Upon initial access, the UI presents a login window and only denies entry if the username already exists. Additionally, the interface transmits user information and messages to the backend server while also broadcasting MUCA’s messages received from the backend server, ensuring they are visible for all participants.
@@
The backend server operates on a locally hosted machine for experimentation purposes. It leverages WebSocket protocol for bi-directional communications to enable multi-user conversations. The backend server is responsible for monitoring incoming messages from all users, and distributing these messages out to the other users. It also manages broadcasting system messages and processing login verification requests. The backend server maintains a record of all connected users, including MUCA, which is a special user that also communicates with other users via the backend.
@@
MUCA establishes an asynchronous connection with the backend server. Upon receiving incoming messages from users, the MUCA determines the appropriate dialog acts, taking chat history and other relevant factors into consideration (as elaborated in Section 3). Subsequently, MUCA’s response is sent back to the backend server for broadcasting to all users. Incoming messages are accumulated in a queue and are processed in batches to prevent an excessive number of rapid API calls.
A.2.6 More Experimental Results
@@
Chit-chat Group Conversation using MUCA: Figure 19 shows the example about chit-chat on topic: During today’s chat, each attendee is going to share their personal choices of prioritizing compensation and work-life-balance when seeking a new job. Hint: encourage people to share their personal choices. Role: a group of attendees who are seeking new jobs.
@@
For chit-chat topics (non-goal-oriented communication), we found that MUCA does not play the same important roles as in goal-oriented communications, since the goal in chit-chat is sharing opinions rather than reaching agreement. In this context, summarizing, voting, or similar functionalities are no longer essential.

### 16 Feb 2024 v3

@@ A.3 User Simulation
Figure 17 shows the prompt templates where purple texts represent placeholders and will be replaced when generating the input prompt for inference. All prompts use a CoT style. For example, a virtual user ID is first determined for the next turn and then its speaking role is decided based on the user ID. Figure 18 shows two examples of MUS for different topics. Examples demonstrate that MUS is able to mimic different user’s behavior based on their speaking roles and utterance traits.
A.4 System Design
@@
The user interface (UI), designed with JavaScript, HTML, and CSS, is a static single-page web application that is responsible for managing user login and facilitating communication with the backend server. Upon initial access, the UI presents a login window and only denies entry if the username already exists. Additionally, the interface transmits user information and messages to the backend server while also broadcasting MUCA’s messages received from the backend server, ensuring they are visible to all participants.
@@
The backend server operates on a locally hosted machine for experimentation purposes. It leverages WebSocket protocol for bi-directional communications to enable multi-user conversations. The backend server is responsible for monitoring incoming messages from all users, and distributing these messages out to the other users. It also manages broadcasting system messages and processes login requests. The backend server maintains a record of all connected users, including MUCA, which is a special user that also communicates with other users via the backend.
@@
MUCA establishes an asynchronous connection with the backend server. Upon receiving incoming messages from users, MUCA determines appropriate dialog acts, taking chat history and relevant factors into consideration (as elaborated in Sec. 3). Subsequently, MUCA’s response is sent back to the backend server for broadcasting. Incoming messages are accumulated in a queue and are processed periodically to better handle API call burst.
A.5 Baseline-small Prompting
Figure 19 shows the prompting of baseline-small. Given the input of user-input information $I$, short-term context window $U_{N_{sw},i}$, and participants’ names, baseline-small will generate the response based on the instructions. baseline-small is also equipped with the capabilities of keeping silent and direct chatting. Examples and comparison are given in Appendix 4.2.2.
A.6 More Experimental Results
Chit-chat Group Conversation using MUCA: Figure 20 shows the example for a chit-chat topic:
@@ A.6 More Experimental Results
  • Topic: During today’s chat, each attendee is going to share their personal choices of prioritizing compensation and work-life-balance when seeking a new job. Hint: encourage people to share their personal choices. Participant roles: they are seeking new jobs.
@@
For chit-chat topics (non-goal-oriented communication), we found that MUCA does not play the same important roles as in goal-oriented communications, since the goal in chit-chat is sharing opinions rather than reaching agreements. In this context, summarizing, voting, or similar functionalities are less critical compared to goal-oriented group conversations.